OpenStudy (anonymous):

Which of the following is equivalent to log 7^ 50 rounded to three decimal places?

hartnn (hartnn):

\(\huge \log x^n=n\log x\)

hartnn (hartnn):

so what will be log 7^50 ??

OpenStudy (anonymous):

log 7^50 = 50 log 7 = 50 x 0.8451 = 42.255

OpenStudy (anonymous):

have you got it now..??

hartnn (hartnn):

@Kashan please do not provide complete solution, let the asker solve it.

hartnn (hartnn):

read: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@Kashan options: 2.010 1.772 0.854 0.497 my bad, there is no ^ sign in the question

hartnn (hartnn):

then use this: \(\huge logxy=logx+logy\)

hartnn (hartnn):

what is the sign between 7 and 50?

OpenStudy (anonymous):

is it 7 multiplies by 50 or 7.50...??

OpenStudy (anonymous):

\[\log_{7} 50\]

OpenStudy (anonymous):

thats what it looks like

hartnn (hartnn):

ohh \(\huge\log_a b=\frac{log b}{log a}\)

OpenStudy (anonymous):

O that is called... log of 50 with base seven.. ok i'll solve it..

hartnn (hartnn):

so log 50/ log 7 =?

OpenStudy (anonymous):

option 1 is right answer..

OpenStudy (anonymous):

thank you guys for all your help!

OpenStudy (anonymous):

no problem.. :)